# Revisiting Locally Differentially Private Protocols: Towards Better Trade-offs in Privacy, Utility, and Attack Resistance

### v2

@@ IV LDP Frequency Estimation Protocols
In this section, we provide a concise overview of eight state-of-the-art LDP frequency estimation protocols. We systematically describe each mechanism based on three key functions: Encoding, Perturbation and Aggregation which collectively define their operation. Additionally, we introduce an Attacking function for each protocol, highlighting potential vulnerabilities to privacy attacks. For protocols in which the expected ASR is not available in the literature [28], we derive it and present the detailed analyses in Appendix -A.
@@ IV-A Generalized Randomized Response (GRR)
Perturbation. The perturbation function of GRR is given by:
@@
Aggregation. In GRR, each output value $i$ supports the corresponding input $i$, resulting in the support set $\mathbb{1}_{\mathrm{GRR}}={y}$. GRR is a pure protocol with $p^{*}=p$ and $q^{*}=q$. The server estimates the frequency using Equation (2), with the following analytical MSE:
@@
Attacking. From Equation (5), it follows that $\Pr[y=x]>\Pr[y=x^{\prime}]$ for all $x^{\prime}\in{\mathcal{X}}\setminus\{x\}$. Thus, the optimal attack strategy for GRR is to predict $\hat{x}={\mathcal{A}}_{\mathrm{GRR}}(y)=y$. The expected ASR for GRR is given by:
$\mathbb{E}[\text{ASR}]_{\mathrm{GRR}}=\frac{e^{\varepsilon}}{e^{\varepsilon}+k-1}\mathrm{.}$ (7)
@@ IV-B Subset Selection (SS)
Aggregation. In SS, each value $i$ in the output subset $\mathbf{y}$ supports the corresponding input value $i$. Thus, the support set for SS is $\mathbb{1}_{\mathrm{SS}}=\{x\mid x\in\mathbf{y}\}$. This protocol is pure, with: $p^{*}=p=\frac{\omega e^{\varepsilon}}{\omega e^{\varepsilon}+k-\omega}$ and $q^{*}=\frac{\omega e^{\varepsilon}(\omega-1)+(k-\omega)\omega}{(k-1)(\omega e^{\varepsilon}+k-\omega)}$. The server estimates the frequency using Equation (2), with the following analytical MSE:
@@
The optimal subset size that minimizes the MSE of SS in Equation (8) is $\omega=\max\left(1,\left\lfloor\frac{k}{e^{\varepsilon}+1}\right\rceil\right)$ [20, 21].
@@
$\mathbb{E}[\text{ASR}]_{\mathrm{SS}}=\frac{e^{\varepsilon}}{\omega e^{\varepsilon}+k-\omega}\mathrm{.}$ (9)
@@ IV-C Unary Encoding (UE)
Encoding. $\texttt{Encode}(x)=[0,\ldots,0,1,0,\ldots,0]$ is a binary vector with a single $1$ at position $x$ and all other positions set to 00.
@@
in which y is sent to the server. There are two variations of UE mechanisms: (i) Symmetric UE (SUE) [3] that selects $p=\frac{e^{\varepsilon/2}}{e^{\varepsilon/2}+1}$ and $q=\frac{1}{e^{\varepsilon/2}+1}$; and (ii) Optimized UE (OUE) [22] that selects $p=\frac{1}{2}$ and $q=\frac{1}{e^{\varepsilon}+1}$ to minimize the MSE in Equation (11) below.
Aggregation. A reported bit vector $\mathbf{y}$ is considered to support an input $i$ if $\mathbf{y}_{i}=1$. Therefore, the support set for UE protocols is defined as $\mathbb{1}_{\mathrm{UE}}=\{i|\textbf{y}_{i}=1\}$. UE protocols are pure with $p^{*}=p$ and $q^{*}=q$. The server estimates the frequency using Equation (2), with the corresponding MSE as:
@@
$\mathbb{E}[\text{ASR}]_{\mathrm{UE}}=$ $\,(1-p)\cdot(1-q)^{k-1}\cdot\frac{1}{k}$ (12)
$+\sum_{m=1}^{k}p\cdot\frac{1}{m}\cdot\binom{k-1}{m-1}q^{m-1}(1-q)^{(k-1)-(m-1)}\mathrm{.}$
We defer the derivation of Equation 12 to Appendix -A1.
@@ IV-D Local Hashing (LH)
LH protocols [22, 5] use hash functions to map the input data $x\in{\mathcal{X}}$ to a new domain $[g]$, and then obfuscates the hash value with GRR. Let $\mathscr{H}$ be a universal hash function family such that each hash function $\mathrm{H}\in\mathscr{H}$ hashes a value $x\in{\mathcal{X}}$ into $[g]$ (i.e., $\mathrm{H}:[k]\rightarrow[g]$).
Encoding. $\texttt{Encode}(x)=\langle\mathrm{H},h\rangle$, in which $\mathrm{H}\in\mathscr{H}$ is chosen uniformly at random, and $h=\mathrm{H}(x)$. There are two variations of LH mechanisms: (i) Binary LH (BLH) [5] that just sets $g=2$, and (ii) Optimized LH (OLH) [22] that selects $g=\lfloor e^{\varepsilon}+1\rceil$ to minimize the MSE in Equation (14) below.
@@
Aggregation. For each reported tuple $\langle\mathrm{H},y\rangle$, the support set for LH protocols consists of all values $x\in{\mathcal{X}}$ that hash to $y$, denoted as $\mathbb{1}_{\mathrm{LH}}=\{x\mid\mathrm{H}(x)=y\}$. LH protocols are pure with $p^{*}=p$ and $q^{*}=\frac{1}{g}$. The server estimates the frequency using Equation (2) with the following analytical MSE:
@@
$\mathbb{E}[\text{ASR}]_{\mathrm{LH}}=\frac{e^{\varepsilon}}{(e^{\varepsilon}+g-1)\cdot\max\left\{\frac{k}{g},1\right\}}\mathrm{.}$ (15)
@@ IV-E1 Summation with HE (SHE)
Aggregation. Since Laplace noise with mean 00 is added to each vector independently, the server estimates the frequency using the sum of the noisy reports: $\hat{\mathbf{f}}=\left\{\sum_{j=1}^{n}\mathbf{y}^{j}_{i}\right\}_{i\in[k]}$. This aggregation method for SHE does not provide a support set and is not pure. The analytical MSE of this estimation is:
@@
$\mathbb{E}[\text{ASR}]_{\mathrm{SHE}}=\Pr\left[y_{x}>\max_{i\neq x}y_{i}\right].$ (17)
We defer the derivation of Equation (17) to Appendix -A2.
@@ IV-E2 Thresholding with HE (THE)
Aggregation. A reported bit vector $\mathbf{y}$ is viewed as supporting an input $i$ if $\mathbf{y}_{i}>\theta$. Therefore, the support set for THE is $\mathbb{1}_{\mathrm{THE}}=\{i\hskip 2.84544pt|\hskip 2.84544pt\textbf{y}_{i}\hskip 2.84544pt>\hskip 2.84544pt\theta\}$. The THE mechanism is pure with $p^{*}=p$ and $q^{*}=q$. The server estimates the frequency using Equation (2) with the following analytical MSE:
@@
The optimal threshold value that minimizes the protocol’s MSE in Equation (20) is within $\theta\in(0.5,1)$ [22].
@@
$\mathbb{E}[\mathrm{ASR}]_{\mathrm{THE}}=(1-p)(1-q)^{k-1}\cdot\frac{1}{k}+\frac{p}{kq}\left(1-(1-q)^{k}\right)\mathrm{.}$ (21)
We defer the derivation of Equation (21) to Appendix -A3.
V Refining LDP Protocols
The existing literature has traditionally proposed mechanisms like SS, OUE, OLH and THE, focusing solely on minimizing MSE for a given privacy budget $\varepsilon$. However, these protocols exhibit varying vulnerabilities to privacy and security attacks under the same $\varepsilon$ value, leaving room for improvement in both privacy guarantees and estimation accuracy. To address this, we introduce a general multi-objective optimization framework for refining LDP frequency estimation protocols, enabling adaptive parameter tuning based on multiple privacy and utility considerations.
V-A Multi-Objective Optimization Framework
We frame the protocol refinement as a general multi-objective optimization problem, in which the goal is to jointly minimize a set of metrics that characterize distinct desirable properties:
$\min_{\theta\in\Theta}\quad\left(\mathcal{P}(\theta),\;\mathcal{U}(\theta),\;\mathcal{S}(\theta),\;\mathcal{C}(\theta)\right)\mathrm{,inwhich}$ (22)
  • $\mathcal{P}$ quantifies the LDP protocol’s vulnerability to privacy attacks, such as reconstruction [28, 29], inference [27, 9, 33] or re-identification [26, 25];
  • $\mathcal{U}$ reflects utility degradation, often via MSE, Mean Absolute Error (MAE) [19], or Fisher Information [32];
  • $\mathcal{S}$ captures security robustness to adversarial manipulation [36] or data poisoning [34];
  • $\mathcal{C}$ measures communication cost, e.g., bandwidth or message size [23].
Each dimension may be relevant depending on the application context, and the goal is to explore Pareto-efficient solutions that avoids sacrificing one objective disproportionately for gains in another.
@@ V-B Specialization to Data Reconstruction and MSE Trade-off
While the full multi-objective framework in Section V-A is general, in this paper we focus on a widely applicable and analytically tractable two-objective setting. Specifically, we optimize the trade-off between:
  • Privacy attacks, measured via the ASR under data reconstruction, which is the most fundamental and consequential threat in privacy-preserving data analysis [37, 38, 39]. Their severity lies in the fact that once an adversary successfully reconstructs individual-level data, a wide range of downstream privacy threats, such as re-identification, attribute inference, profiling or membership inference, can be carried out with significantly higher success rates. As such, reconstruction risk serves as a powerful proxy for broader privacy leakage, making it a natural privacy criterion and more interpretable than $\varepsilon$.
@@
  • Utility, measured as the variance (i.e., MSE) in frequency estimation, which directly quantifies the deviation between true and estimated statistics and is central in frequency estimation tasks [30, 22, 31]. Moreover, minimizing MSE inherently reduces other utility losses such as MAE and Fisher Information loss [32] through established theoretical relationships like the Central Limit Theorem and Cramér-Rao bounds [40, 41, 42].
@@
This ASR-MSE formulation captures a critical privacy-utility frontier that is central to many LDP deployments, and both metrics admit closed-form expressions for a wide range of protocols (see Section IV).
V-C Protocol Selection via Pareto Optimization
Given that ASR and MSE represent competing objectives, we search for Pareto-efficient configurations of protocol parameters. To this end, we first evaluate the objective metrics across a predefined parameter grid [43] to construct the Pareto frontier. Then, to select a concrete operating point on this frontier, we consider several established multi-objective selection strategies [44]:
  • Utopia Point Minimization, which picks the solution closest to the ideal point ($\mathrm{ASR}=0$, $\mathrm{MSE}=0$);
  • Weighted Scalarization, which minimizes a linear combination of ASR and MSE via tunable weights;
  • Elbow (Knee) Method, which identifies the point of maximum curvature on the frontier;
  • $\boldsymbol{\epsilon}$-Constraint Selection, which filters solutions with ASR $\leq\epsilon$ and then picks the one with minimum MSE (we default it back to utopia if no feasible point exists);
  • Hypervolume Contribution, which selects the Pareto point contributing the largest area (in 2D) to the dominated space relative to a reference point (e.g., (1,1)), thereby improving coverage of the frontier;
  • Augmented Tchebycheff Scalarization, which selects the point minimizing $\max_{i}\{w_{i}f_{i}\}+\rho\sum_{i}w_{i}f_{i}$, emphasizing worst-case performance while regularizing toward balanced objectives.
Among these, weighted scalarization stands out as a practical default due to its conceptual simplicity, interpretability and tunable nature. Empirically, we observed that all selection methods, except for the Elbow and $\boldsymbol{\epsilon}$-Constraint methods, tend to identify similar parameters (see Figure 9 in Appendix -B). Therefore, while our framework and implementation support multiple optimization criteria, we adopt weighted scalarization as the default approach for selecting protocol parameters:
$\underset{\mathbf{\Theta}}{\mathrm{min}}$ $J(\mathbf{\Theta})=w_{\text{ASR}}\cdot\mathbb{E}[\text{ASR}]+w_{\mathrm{MSE}}\cdot\mathrm{MSE}\mathrm{,}$ (23)
@@ V-C Protocol Selection via Pareto Optimization
The weight $w_{\text{ASR}}$ and $w_{\mathrm{MSE}}$ in Equation (23) can be tuned depending on the context. This flexibility enables LDP deployments to be tailored to different risk-utility trade-offs, enhancing applicability across diverse real-world scenarios. For privacy-critical applications, increasing $w_{\text{ASR}}$ yields safer mechanisms while for analytics-heavy settings, higher $w_{\text{MSE}}$ emphasizes utility. Importantly, setting $w_{\text{MSE}}=1$ recovers the original protocols, showing that our refined mechanisms generalize rather than replace their classical counterparts.
V-D Adaptive Parameter Optimization for LDP Protocols
[…]
V-D1 Adaptive Subset Selection (ASS)

### v3

@@ IV Overview of Pure LDP Frequency
In this section, we provide a concise overview of eight state-of-the-art pure LDP frequency estimation protocols, selected as representative mechanisms underlying widely deployed industrial systems (e.g., SUE-based in Microsoft [5] telemetry, SS in Google Gboard [39] and OLH-based in Apple [4]). We systematically describe each mechanism based on three key functions: Encoding, Perturbation and Aggregation, which collectively define their operation. Additionally, we introduce an Attacking function for each protocol, highlighting potential vulnerabilities to privacy attacks. For protocols in which the expected ASR is not available in the literature [30], we derive it and present the detailed analyses in Appendix -A.
@@ IV-A Generalized Randomized Response (GRR)
Perturbation. The perturbation function of GRR is:
@@
Aggregation. In GRR, each output value $i$ supports the corresponding input $i$, resulting in the support set $\mathbb{1}_{\mathrm{GRR}}={y}$. GRR is a pure protocol with $p^{*}=p$ and $q^{*}=q$. The server estimates the frequency using Eq. (2), with the following analytical MSE:
@@
Attacking. From Eq. (9), it follows that $\Pr[y=x]>\Pr[y=x^{\prime}]$ for all $x^{\prime}\in{\mathcal{X}}\setminus\{x\}$. Thus, the optimal attack strategy for GRR is to predict $\hat{x}={\mathcal{A}}_{\mathrm{GRR}}(y)=y$. The expected ASR for GRR is given by:
$\text{ASR}_{\mathrm{GRR}}=\frac{e^{\varepsilon}}{e^{\varepsilon}+k-1}\mathrm{.}$ (11)
@@ IV-B Subset Selection (SS)
Aggregation. In SS, each value $i$ in the output subset $\mathbf{y}$ supports the corresponding input value $i$. Thus, the support set for SS is $\mathbb{1}_{\mathrm{SS}}=\{x\mid x\in\mathbf{y}\}$. This protocol is pure, with: $p^{*}=p=\frac{\omega e^{\varepsilon}}{\omega e^{\varepsilon}+k-\omega}$ and $q^{*}=\frac{\omega e^{\varepsilon}(\omega-1)+(k-\omega)\omega}{(k-1)(\omega e^{\varepsilon}+k-\omega)}$. The server estimates the frequency using Eq. (2), with the following analytical MSE:
@@
The optimal subset size that minimizes the MSE of SS in Eq. (12) is $\omega=\max\left(1,\left\lfloor\frac{k}{e^{\varepsilon}+1}\right\rceil\right)$ [21, 22].
@@
$\text{ASR}_{\mathrm{SS}}=\frac{e^{\varepsilon}}{\omega e^{\varepsilon}+k-\omega}\mathrm{.}$ (13)
@@ IV-C Unary Encoding (UE)
Encoding. $\texttt{Encode}(x)=[0,\ldots,0,1,0,\ldots,0]$ is a binary vector with a single $1$ at position $x$ and all other positions set to 0.
@@
in which y is sent to the server. There are two variations of UE mechanisms: (i) Symmetric UE (SUE) [3] that selects $p=\frac{e^{\varepsilon/2}}{e^{\varepsilon/2}+1}$ and $q=\frac{1}{e^{\varepsilon/2}+1}$; and (ii) Optimized UE (OUE) [23] that selects $p=\frac{1}{2}$ and $q=\frac{1}{e^{\varepsilon}+1}$ to minimize the MSE in Eq. (15) below.
Aggregation. A reported bit vector $\mathbf{y}$ is considered to support an input $i$ if $\mathbf{y}_{i}=1$. Therefore, the support set for UE protocols is defined as $\mathbb{1}_{\mathrm{UE}}=\{i|\textbf{y}_{i}=1\}$. UE protocols are pure with $p^{*}=p$ and $q^{*}=q$. The server estimates the frequency using Eq. (2), with the corresponding MSE as:
@@
$\mathrm{ASR}_{\mathrm{UE}}=\frac{1}{k}\left[(1-p)(1-q)^{k-1}+\frac{p}{q}\left(1-(1-q)^{k}\right)\right]$ (16)
We defer the derivation of Eq. (16) to Appendix -A1.
@@ IV-D Local Hashing (LH)
LH protocols [23, 6] use hash functions to map the input data $x\in{\mathcal{X}}$ to a new domain $[g]$, and then obfuscate the hash value with GRR. Let $\mathscr{H}$ be a universal hash function family such that each hash function $\mathrm{H}\in\mathscr{H}$ hashes a value $x\in{\mathcal{X}}$ into $[g]$ (i.e., $\mathrm{H}:[k]\rightarrow[g]$).
Encoding. $\texttt{Encode}(x)=\langle\mathrm{H},h\rangle$, in which $\mathrm{H}\in\mathscr{H}$ is chosen uniformly at random, and $h=\mathrm{H}(x)$. There are two variations of LH mechanisms: (i) Binary LH (BLH) [6] that just sets $g=2$, and (ii) Optimized LH (OLH) [23] that selects $g=\lfloor e^{\varepsilon}+1\rceil$ to minimize the MSE in Eq. (18) below.
@@
Aggregation. For each reported tuple $\langle\mathrm{H},y\rangle$, the support set for LH protocols consists of all values $x\in{\mathcal{X}}$ that hash to $y$, denoted as $\mathbb{1}_{\mathrm{LH}}=\{x\mid\mathrm{H}(x)=y\}$. LH protocols are pure with $p^{*}=p$ and $q^{*}=\frac{1}{g}$. The server estimates the frequency using Eq. (2) with the following analytical MSE:
@@
$\text{ASR}_{\mathrm{LH}}=\frac{e^{\varepsilon}}{(e^{\varepsilon}+g-1)\cdot\max\left\{\frac{k}{g},1\right\}}\mathrm{.}$ (19)
@@ IV-E1 Summation with HE (SHE)
Aggregation. Since Laplace noise with mean 0 is added to each vector independently, the server estimates the frequency using the sum of the noisy reports: $\hat{\mathbf{f}}=\left\{\sum_{j=1}^{n}\mathbf{y}^{j}_{i}\right\}_{i\in[k]}$. This aggregation method for SHE does not provide a support set and is not pure. The analytical MSE of this estimation is:
@@
$\text{ASR}_{\mathrm{SHE}}=\Pr\left[y_{x}>\max_{i\neq x}y_{i}\right].$ (21)
We defer the derivation of Eq. (21) to Appendix -A2.
@@ IV-E2 Thresholding with HE (THE)
Aggregation. A reported bit vector $\mathbf{y}$ is viewed as supporting an input $i$ if $\mathbf{y}_{i}>\theta$. Therefore, the support set for THE is $\mathbb{1}_{\mathrm{THE}}=\{i\hskip 2.84544pt|\hskip 2.84544pt\textbf{y}_{i}\hskip 2.84544pt>\hskip 2.84544pt\theta\}$. The THE mechanism is pure with $p^{*}=p$ and $q^{*}=q$. The server estimates the frequency using Eq. (2) with the following analytical MSE:
@@
The optimal threshold value that minimizes the protocol’s MSE in Eq. (24) is within $\theta\in(0.5,1)$ [23].
@@
$\mathrm{ASR}_{\mathrm{THE}}=(1-p)(1-q)^{k-1}\cdot\frac{1}{k}+\frac{p}{kq}\left(1-(1-q)^{k}\right)\mathrm{.}$ (25)
We defer the derivation of Eq. (25) to Appendix -A3.
V Attack-Aware Evaluation Framework for LDP
The existing literature has traditionally proposed mechanisms like SS, OUE, OLH and THE, with the primary objective of minimizing estimation error for a given privacy budget $\varepsilon$. However, protocols offering identical $\varepsilon$-LDP guarantees can exhibit substantially different vulnerabilities to inference and poisoning attacks, leading to markedly different real-world privacy and integrity risks. Motivated by this observation, we propose an attack-aware evaluation framework for LDP frequency estimation protocols, designed to systematically analyze how estimation utility interacts with adversarial vulnerability under different threat models and protocol configurations.
V-A General Evaluation Framework
We evaluate LDP protocols through a unified attack-aware instantiation, in which each mechanism is characterized by its behavior along two complementary axes beyond $\varepsilon$, capturing key system desiderata:
$\left(\mathcal{A}(\theta),\;\mathcal{U}(\theta)\right),$ (26)
in which $\theta\in\Theta$ denotes the protocol parameters (e.g., $\varepsilon$ or design choices), $\mathcal{U}$ measures utility loss and $\mathcal{A}$ quantifies the protocol attackability under a specified adversarial model. Specifically, the attackability component $\mathcal{A}$ can be instantiated to capture different classes of threats:
  • Privacy attackability ($\mathcal{A}_{\mathrm{priv}}$), including data reconstruction [30, 31], attribute inference [29, 10, 35] or re-identification [28, 27];
  • Integrity attackability ($\mathcal{A}_{\mathrm{sec}}$), such as vulnerability to data poisoning [34] or adversarial manipulation [41].
Utility loss $\mathcal{U}$ is typically quantified using metrics based on estimation error, such as MSE, Mean Absolute Error (MAE) [20], or Fisher Information [42].
This formulation provides a principled lens for reasoning about the inherent trade-offs between estimation accuracy and adversarial vulnerability in the design of LDP protocols.
@@ V-B Specialization to Data Reconstruction and MSE Trade-off
In this work, we focus on a widely applicable and analytically tractable instantiation that captures the fundamental privacy-utility tension in LDP. Specifically, we consider:
  • Privacy attackability, measured via the ASR under data reconstruction attacks in Eq. (8). Reconstruction attacks directly target users’ private values and admit closed-form analyses for all eight state-of-the-art LDP protocols considered in this paper. Moreover, successful reconstruction can serve as a proxy for other downstream privacy risks [36, 37], such as profiling and re-identification (see Example 1) as well as attribute inference (see Example 2), making ASR a natural and interpretable indicator of broader privacy leakage.
@@
  • Utility, measured by the variance (MSE) of frequency estimation, which directly quantifies estimation accuracy and is central to frequency estimation tasks [32, 23, 33]. Minimizing MSE also implicitly controls other utility losses such as MAE and Fisher Information through established theoretical relationships like the Central Limit Theorem and Cramér-Rao bounds [43, 44, 45].
@@
This ASR-MSE instantiation yields a meaningful privacy-utility frontier that is central to many practical LDP deployments, and leverages closed-form expressions available for a wide range of protocols (see Section IV).
Protocol Selection via Pareto Optimization. Since ASR and MSE represent inherently competing objectives, each LDP protocol can be characterized by a set of Pareto-efficient configurations that balance privacy leakage and utility. To identify these configurations, we evaluate both metrics across a predefined parameter grid [46], forming the analytical Pareto frontier that captures the achievable trade-offs for each mechanism. Once the Pareto frontier is obtained, we select representative operating points using established multi-objective selection criteria (e.g., Utopia Point, Weighted Scalarization, Elbow “Knee” Method, etc) [47]. Among these, weighted scalarization stands out as a practical default due to its conceptual simplicity, interpretability and tunable nature. Empirically, we observed that all selection methods, except for the Elbow and $\boldsymbol{\epsilon}$-Constraint methods, tend to identify similar parameters (see Figure 7 in Appendix -B). For clarity and consistency, we therefore report results using the scalarized formulation:
$\underset{\mathbf{\Theta}}{\mathrm{min}}$ $J(\mathbf{\Theta})=w_{\text{ASR}}\cdot\text{ASR}+w_{\mathrm{MSE}}\cdot\mathrm{MSE}\mathrm{,}$ (27)
@@
Varying the weights $(w_{\text{ASR}},w_{\mathrm{MSE}})$ enables controlled exploration of privacy-utility preferences: higher $w_{\text{ASR}}$ emphasizes stronger protection, while larger $w_{\mathrm{MSE}}$ favors utility. This parameterization provides a unified way to visualize trade-offs across protocols and privacy regimes. Importantly, setting $w_{\text{MSE}}=1$ recovers the original protocols, showing that our refined mechanisms generalize rather than replace their classical counterparts.
Remark 1 (Attack-Agnostic Nature of the Framework).
The proposed attack-aware framework in Section V-A is not restricted to reconstruction attacks. More precisely, it is attack-agnostic and can accommodate arbitrary adversarial models, provided that the corresponding attackability metric can be evaluated, either analytically or via empirical (Monte Carlo) estimation.
In Section VI-E, we illustrate this flexibility by incorporating integrity-oriented attacks through data poisoning attacks [34]. While different attack metrics may interact with protocol parameters in distinct ways, the framework naturally supports their joint analysis, either via complementary two-dimensional trade-offs or through higher-dimensional visualizations and constrained selection strategies.
V-C Refining LDP Protocols
Using our two-objective framework defined in Eq. (27), we extend four state-of-the-art LDP protocols to introduce adaptive counterparts. Each adaptive protocol selects an optimal parameter $\mathbf{\Theta}$ to achieve a better trade-off between ASR and MSE, rather than focusing solely on utility. The optimization process varies across protocols, adapting their internal parameters to balance privacy protection and estimation accuracy. Importantly, our reparametrization preserves the original $\varepsilon$-LDP guarantee by design. For instance, the subset size $\omega$ in ASS and the probability pair $(p,q)$ in AUE are explicitly constrained to satisfy the $\varepsilon$-LDP guarantee; ALH inherits the privacy guarantee of GRR applied to a hashed domain; and ATHE applies a post-processing threshold, which does not affect the privacy guarantee. We now describe each adaptive mechanism in detail.
V-C1 Adaptive Subset Selection (ASS)